# MMGRec: Multimodal Generative Recommendation with Transformer Model

### v1

@@ 4.1. Datasets & Evaluation Metrics
For each dataset, we randomly select 80% of the historical interactions per user to constitute the training set, 10% for validation, and the remaining 10% for testing. The validation and testing sets are used for hyper-parameter tuning and performance evaluation, respectively. We adopt Recall (R@$K$) and Normalized Discounted Cumulative Gain (N@$K$) as evaluation metrics for top-$K$ recommendation and preference ranking. Defaulting $K=10$, we report the average metric value for all users in the testing set.
@@ 4.2. Baselines
To ensure consistency, we adopt the publicly available implementations of the baselines. Note that we did not select the existing generative recommendation model TIGER (Rajput et al., 2023) as baseline, since it only works with sequential recommendation rather than our task.
@@ 5.1. Performance Comparison (RQ1)
  • MMGRec consistently achieves the best performance in all cases. In particular, improvements over the strongest competitor w.r.t. NDCG@10 are 7.17%, 6.79%, and 6.58% in MovieLens, TikTok, and Kwai, respectively. Additionally, we conduct one-sample t-tests, which reveal that the improvements of MMGRec are statistically significant (p-value ¡ 0.05). The primary difference between MMGRec and baselines lies in whether the recommendation paradigm is generative or traditional. Hence, we attribute the significant improvements to introducing a Transformer-based generative paradigm in multimodal recommendation.
@@
R@10 N@10 R@10 N@10 R@10 N@10
@@ 5.2. Ablation Study (RQ2)
R@10 N@10 R@10 N@10 R@10 N@10
@@ 5.2.2. Effect of Graph RQ-VAE
Methods MovieLens TikTok Kwai
R@10 N@10 R@10 N@10 R@10 N@10
HK-Means 0.2520 0.1683 0.1137 0.0742 0.0469 0.0402
Random 0.2762 0.1872 0.1199 0.0785 0.0544 0.0452
Ours 0.2804 0.1898 0.1269 0.0802 0.0567 0.0470
@@ 5.2.3. Effect of Relation-Aware Self-Attention
R@10 N@10 R@10 N@10 R@10 N@10
@@ 5.4. Efficiency Study (RQ4)
Instead of computing and ranking similarity, MMGRec employs autoregressive decoding to directly generate Rec-IDs identifying items during inference. To assess the efficiency of the two inference approaches, we conduct comparative experiments using identical computing configurations and record their inference time. Figure 5 presents an efficiency comparison between MMGRec and MF regarding average inference time per user across different-scale Kwai datasets. 1/8 Kwai denotes randomly selecting one in eight items from the complete dataset. The main observations are as follows:
@@
In scenarios where the item count is low (e.g., $1/16\sim 1/4$ Kwai), MMGRec exhibits inefficiency, primarily due to the basic computational demands of autoregressive decoding. However, as the number of items increases (e.g., $1/4\sim 1$ Kwai), the inference time of MF experiences a significant rise. This escalation is attributed to the fact that item count linearly impacts the time complexity of inner product computation and sorting. In contrast, MMGRec demonstrates enhanced efficiency as its inference time remains relatively stable. This stability arises from the unchanged number of autoregressive decoding steps, even with an increase in the number of items. These findings suggest that the generative paradigm holds the potential to deliver more efficient inference for large-scale recommendation.
@@
Figure 5. Inference time on different-scale Kwai datasets. The unit of time is milliseconds (ms).
6. Related Work
6.1. Multimodal Recommendation
Research on multimodal recommendation aims to enhance recommender systems by leveraging multimodal information. Existing methodologies typically employ a combination of CF and content-based recommendation techniques (Liu et al., 2023). Early studies aim to integrate multimodal content for more rational representation learning. For example, the pioneering work VBPR (He and McAuley, 2016) maps visual information to the representation space and concatenates it with CF representations. Subsequent work like ACF (Chen et al., 2017) and UVCAN (Liu et al., 2019) adaptively select multimodal features using different levels of attention mechanisms to learn representations of users and items. In recent years, the research trend turns to exploring the relationship between modal information and users/items. For instance, the representative work MMGCN (Wei et al., 2019) introduces a modality-aware GCN into multimodal recommendation tasks to aggregate and propagate multimodal information on the user-item graph. HHFAN (Cai et al., 2021) utilizes a modality-aware heterogeneous information graph to explore more complex relationships among users, micro-videos, and multimodal information, thereby improving recommendation performance.
6.2. Transformer-Based Recommendation Model

### v2

@@ 5.1. Datasets & Evaluation Metrics
For each dataset, we randomly select 80% of the historical interactions per user to constitute the training set, 10% for validation, and the remaining 10% for testing. The validation and testing sets are used for hyper-parameter tuning and performance evaluation, respectively. We adopt Recall (Recall@$K$) and Normalized Discounted Cumulative Gain (NDCG@$K$) as evaluation metrics for top-$K$ recommendation and preference ranking. Defaulting $K=10$, we report the average metric value for all users in the testing set.
@@ 5.2. Baselines
To ensure consistency, we adopt the publicly available implementations of the baselines. Note that we do not include TIGER (Rajput et al., 2023) as a baseline. TIGER is designed for sequential recommendation and requires temporally ordered user-item interactions. In contrast, our work focuses on non-sequential collaborative filtering, and the datasets used in our experiments do not contain temporal labels. Moreover, TIGER constructs semantic IDs solely from textual features, while our method relies on multimodal item representations. Therefore, TIGER is not directly applicable to our setting.
@@ 6. Experimental Results
  • RQ5: How do Rec-ID design choices affect collisions and recommendation performance?
@@ 6.1. Performance Comparison (RQ1)
  • MMGRec consistently achieves the best performance in all cases. In particular, improvements over the strongest competitor w.r.t. NDCG@10 are 7.17%, 6.79%, and 6.58% in MovieLens, TikTok, and Kwai, respectively. Additionally, we conduct one-sample t-tests, which reveal that the improvements of MMGRec are statistically significant (p-value << 0.05). The primary difference between MMGRec and baselines lies in whether the recommendation paradigm is generative or traditional. Hence, we attribute the significant improvements to introducing a Transformer-based generative paradigm in multimodal recommendation.
@@
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
@@ 6.2. Ablation Study (RQ2)
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
@@ 6.2.1. Effect of Item Representation
Methods MovieLens TikTok Kwai
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
HK-Means 0.2520 0.1683 0.1137 0.0742 0.0469 0.0402
Random 0.2762 0.1872 0.1199 0.0785 0.0544 0.0452
Ours 0.2804 0.1898 0.1269 0.0802 0.0567 0.0470
@@ 6.2.3. Effect of Relation-Aware Self-Attention
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
@@ 6.4. Efficiency Study (RQ4)
Instead of computing and ranking similarity, MMGRec employs autoregressive decoding to directly generate Rec-IDs identifying items during inference. To assess the efficiency of the two inference approaches, we conduct comparative experiments using identical computing configurations and record their inference time. Figure 5 presents a pairwise efficiency comparison between MMGRec and each of MF, LightGCN, and LightGT, in terms of the average inference time per user across Kwai datasets of varying scales. 1/8 Kwai denotes randomly selecting one in eight items from the complete dataset. The main observations are as follows:
@@
For LightGCN and LightGT, the propagated embeddings are precomputed offline, such that the online inference is effectively reduced to inner-product retrieval. Under this setting, the computational complexity of LightGCN is comparable to that of MF, while LightGT exhibits higher complexity due to inner-product computations across multiple modalities. Consequently, our efficiency analysis primarily focuses on comparing MMGRec with MF, which represents the most efficient baseline. In scenarios where the item count is low (e.g., $1/16\sim 1/4$ Kwai), MMGRec exhibits inefficiency, primarily due to the basic computational demands of autoregressive decoding. However, as the number of items increases (e.g., $1/4\sim 1$ Kwai), the inference time of MF experiences a significant rise. This escalation is attributed to the fact that item count linearly impacts the time complexity of inner product computation and sorting. In contrast, MMGRec demonstrates enhanced efficiency as its inference time remains relatively stable. This stability arises from the unchanged number of autoregressive decoding steps, even with an increase in the number of items. These findings suggest that the generative paradigm holds the potential to deliver more efficient inference for large-scale recommendation.
@@
Figure 5. Inference time on different-scale Kwai dataset. The unit of time is milliseconds (ms).
6.5. Analysis of Rec-ID Design (RQ5)
This subsection provides a quantitative analysis of the Rec-ID design, focusing on the effect of the popularity token in mitigating Rec-ID collisions, as well as the impact of Rec-ID length and codebook size on model performance. Due to consistent trends across datasets and space limitations, we report results only on the Kwai dataset, as shown in Figure 6.
Impact of Rec-ID Collision. As shown in Figure 6(a), the bar plots show the collision rate on the Kwai dataset under different data scales, measured as the proportion of items assigned non-unique Rec-IDs. The x-axis value 1/16 indicates a subset sampled from one-sixteenth of the items and their interacting users. As the dataset scale increases, the collision rate rises, and model performance degrades slightly, indicating that Rec-ID collisions have a non-negligible impact in large-scale recommendation scenarios. The line plots compare our popularity token with a random token (as in TIGER). With higher collision rates, the popularity token achieves larger performance gains. This indicates that popularity-aware semantic ordering better distinguishes colliding Rec-IDs and reduces the performance degradation caused by collisions.
Effect of Rec-ID Length. Figure 6(b) shows the performance trends in Recall@10 and NDCG@10 as the Rec-ID length $M$ increases from 2 to 5. To cover the full item set, a shorter Rec-ID results in a larger per-step decoding search space. As $M$ increases from 2 to 4, model performance steadily improves, since overly short Rec-IDs increase decoding difficulty due to a large set of candidate tokens at each step. When $M=4$, the model achieves a good balance between the number of decoding steps and the per-step search space, leading to the best performance. However, when $M$ further increases to 5, more autoregressive steps introduce higher generation difficulty and inference latency, resulting in performance degradation. Therefore, we set $M=4$ by default.
Effect of Codebook Size. To evaluate the impact of codebook size $L$, we fix the Rec-ID length to $M=4$ and vary $L$ over $\{64,128,256,512\}$. As shown in Figure 6(c), performance improves as $L$ increases from 64 to 128, since a small codebook has limited representational capacity and cannot adequately distinguish Rec-ID semantics. When $L=128$, the codebook is sufficiently expressive and achieves the best performance. However, further increasing $L$ rapidly enlarges the decoding search space, forcing the model to select from more candidate tokens at each step, which increases generation uncertainty and degrades performance. These results suggest that codebook size should balance representational capacity and decoding complexity in generative recommendation.
Figure 6. Quantitative analysis of Rec-ID design on the Kwai dataset. (a) Impact of Rec-ID collision and popularity token. (b) Effect of Rec-ID length. (c) Effect of codebook size.